Find the work done in stretching a wire of 1 sq. mm cross-section and 2 m long through 0.1 mm. Given that the Young’s modulus of the material of the wire is 2.0 × 1011 N/m2. 


1
Expert's answer
2021-03-08T08:31:52-0500

The work is:


"W = \dfrac{1}{2}kx^2"

The spring constant "k" of the wire can be found from the Hook's law:


"F = kx"

where "F" is the elastic force, and "x = 0.1mm = 10^{-4}m" is the elongation of the wire.

On the other hand, the Hook's law states:


"\dfrac{F}{A} = E\dfrac{x}{L}"

where "A = 1mm^2 = 10^{-6}m^2" is the cross-section, "L = 2m" is the length of the wire, and "E = 2\times 10^{11}N/m^2" is the Young’s modulus of the material of the wire.

Equating the force from these two equations, obtain:


"kx = \dfrac{EAx}{L}\\
k = \dfrac{EA}{L}"

Substituting this expression into the formula for the work, obtain:


"W = \dfrac{EAx^2}{2L}\\
W = \dfrac{2\times 10^{11}\cdot 10^{-6}\cdot (10^{-4})^2}{2\cdot 2} = 5\times 10^{-4}J"

Answer. "5\times 10^{-4}J".